OpenStudy (anonymous):

If the domain of the square root function f(x) is x<=7, which statement must be true? 7 is subtracted from the x-term inside the radical. The radical is multiplied by a negative number. 7 is added to the radical term. The x-term inside the radical has a negative coefficient.

OpenStudy (anonymous):

s = ut + ½ at2

OpenStudy (anonymous):

s = ut + 1/2 at2

OpenStudy (anonymous):

do you understand that domain in a different form...

OpenStudy (anonymous):

no, sorry.

OpenStudy (anonymous):

@CavebusterWatcher!

OpenStudy (anonymous):

Are you sure it is not x>=7 because none of those answers work.

OpenStudy (anonymous):

Yeah, I'm sure. See for yourself.

jimthompson5910 (jim_thompson5910):

you could do this trick x <= 7 x-7<=7-7 x-7 <= 0 -1*(x-7) >= -1*0 ... notice the sign flip -x + 7 >= 0 7 - x >= 0 ----------------------------------- or you can do it this way x <= 7 x-x <= 7-x 0 <= 7-x 7-x >= 0

jimthompson5910 (jim_thompson5910):

either way, you go from x <= 7 to 7-x >= 0

OpenStudy (anonymous):

Is it the last option?

OpenStudy (anonymous):

The x has a negative coefficient

OpenStudy (anonymous):

I was considering that.

OpenStudy (anonymous):

Yeah, I think I'll go with that.